Given μ=70.5 and σ=4.6 of normal distribution, find 


а. p(65 < х < 75) 

b. p(72 < x < 80) 

c. p(x > 80) 

d. p(x < 65)
Answer #1

answered by: Hasson Ali
Answer #2

a)

for normal distribution z score =(X-μ)/σx
here mean=       μ= 70.5
std deviation   =σ= 4.6000
probability = P(65 = P(-1.2 0.8365-0.1151= 0.7214

b)

probability = P(72 = P(0.33 0.9808-0.6293= 0.3515

c)

probability = P(X>80) = P(Z>2.065)= 1-P(Z<2.07)= 1-0.9808= 0.0192

d)

probability = P(X<65) = P(Z<-1.2)= 0.1151